Question2 1 pts A 1,500,000 SF plant employs 300 people and produces 251,000 LCD panel per month. The plant is idle, on average, 2 months eaclh year for maintenance, repairs, and retooling. The plant has fixed costs of $4,000,000 per month and variable costs of $90 per LCD panel. What is the total operating cost for the plant per YEAR? Express your answer in $to the nearest $1 million.
Answer #1

Annual fixed cost = 4000000*12 = $48000000

Monthly variable cost = 251000*90 = $22590000

No. of months of production in a year= 10

annual variable cost = 22590000*10 = $225900000

So,

Annual operating cost = 48000000+225900000

Annual operating cost = $273900000 or  $273.9 Million or $274 Million